### v3

# Learning county from pixels: Corn yield prediction with attention-weighted multiple instance learning

Abstract
Remote sensing has become a promising tool in yield prediction. Most prior work using satellite imagery for county-level corn yield prediction spatially aggregates all pixels within a county into a single value, potentially overlooking the detailed information and valuable insights offered by more granular data. This research studies each county at the pixel level and applies multiple instance learning to leverage the detailed information within each county. In addition, our method addresses the “mixed pixel” problem caused by the inconsistent resolution between feature datasets and crop masks, which may introduce noise into the model and make corn yield prediction more difficult. Specifically, the attention mechanism is utilized to automatically assign weights to different pixels, which can mitigate the influence of the “mixed pixel” problem. The experimental results show that the developed model outperforms four other machine learning models over the past five years in the U.S. Corn Belt and demonstrates its best performance in 2022, achieving a coefficient of determination ($R^{2}$) value of 0.84 and a root mean square error ($RMSE$) value of 0.83. This paper demonstrates the advantages of our approach from both spatial and temporal perspectives. Furthermore, through an in-depth study of the relationship between mixed pixels and attention, it is verified that our approach can capture important feature information while filtering out noise from mixed pixels.
articletype:
This work was supported by the United States Department of Agriculture (USDA) National Institute of Food and Agriculture, Agriculture and Food Research Initiative Project under Grant 1028199.
1 Introduction
County-level corn yield prediction in the U.S. holds significant importance due to its central role in the country’s agriculture and economy (Shuai and Basso 2022; Yang et al. 2021; Ma et al. 2021a). Accurate predictions enable better planning for supply chain management, price setting, and policy-making, as well as assisting farmers in the decision-making process to optimize resource use. Due to the advantages of efficiency, easy data acquisition, and wide spatial coverage, remote sensing (RS) technology has been widely applied to yield prediction (Deines et al. 2021; Jin et al. 2019; Battude et al. 2016). RS for yield prediction typically involves two main approaches: process-based method (Bastiaanssen and Ali 2003; Huang et al. 2015; Mishra et al. 2013; Lobell and Burke 2010) and learning-based method (Shuai and Basso 2022; Ma et al. 2021a). The former utilizes mathematical models to represent and simulate the physical and biological processes that determine crop growth and development. The latter predicts yield by learning the complex and non-linear relationships between yield and features extracted from RS data. As a data-driven approach, the learning-based method can quickly process large volumes of data, leading to better generalization on new data, without the need for explicitly modeling the physical processes as in the process-based method (Sakamoto, Gitelson, and Arkebauer 2014; Johnson 2014a; Sakamoto, Gitelson, and Arkebauer 2013). This study focuses on exploring the use of a learning-based method for the county-level corn yield prediction task.
Deep learning (Goodfellow, Bengio, and Courville 2016) models are currently the most popular and effective learning-based models. Many studies have investigated the use of various types of neural networks with RS data for the crop yield prediction task. One of the earliest neural networks, the Multi-Layer Perceptron (MLP) (Haykin 1998), consists of multiple layers of nodes in a directed graph and can model complex data (Khaki and Wang 2019; Bazrafshan et al. 2022; Khalifani et al. 2022). The Convolutional Neural Network (CNN) (LeCun, Bengio et al. 1995) uses a mathematical operation called convolution, which is highly effective in processing images (Nevavuori, Narra, and Lipping 2019; Khaki, Wang, and Archontoulis 2020a; Shahhosseini et al. 2021b). For time series modeling, the Recurrent Neural Network (RNN) (Rumelhart et al. 1985) with its internal memory can capture information about previous steps (Khaki, Wang, and Archontoulis 2020b; Fan et al. 2022; Gong et al. 2021). Long Short-Term Memory (LSTM) networks (Hochreiter and Schmidhuber 1997) are good at modeling temporal sequences and capturing phenological trends over time (Jiang et al. 2020). By combining CNN and LSTM, we can utilize both spatial and temporal information (Sun et al. 2019). Furthermore, ResNet (He et al. 2016) employs residual learning, which helps mitigate the vanishing gradient problem in deep networks and improves feature extraction and accuracy (Sagan et al. 2021). Recently, the Transformer (Vaswani et al. 2017) model has significantly improved the performance in various language tasks and has become the foundation of multi-modal combination Liu et al. (2022a). Various Transformer variants, like the Vision Transformer (ViT) (Dosovitskiy 2020) and Swin Transformer (Swin-T) (Liu et al. 2021), have also become mainstream models in remote sensing image processing, demonstrating the potential and rapid development of deep learning for yield prediction tasks (Lin et al. 2023; Krishnan et al. 2024).
In county-level yield prediction, many previous studies treat a county as a unified entity, aggregating all pixels within the county to a single value (e.g. mean value) (Shahhosseini et al. 2021a; Wang et al. 2020; Sakamoto, Gitelson, and Arkebauer 2014; Ma et al. 2021a). This method not only simplifies the model by considerably reducing the computational load but also results in accurate predictions. For example, Wang et al. (2020) develops a dual-branch deep learning framework for county-level winter wheat yield prediction, where each county’s RS information is spatially compressed into the average value. Ma et al. (2021a) employs a Bayesian neural network for both county-level corn yield prediction and its corresponding predictive uncertainty. This study utilizes the median value instead of average to effectively reduce the influence of outliers on the yields (Shahhosseini et al. 2021a). However, aggregating all the pixels within the county to a single value may compromise the integrity of each county’s data, potentially losing detailed information. As a result, missing crucial information could substantially reduce the model’s predictive performance. Therefore, we collect all the pixels of cornfields from each county and utilize them as completely as possible to predict the county’s yield, thereby enhancing the accuracy and reliability of the predictions.
To better utilize the information from pixels within each county, we employ Multiple Instance Learning (MIL) for modeling. MIL (Dietterich, Lathrop, and Lozano-Pérez 1997; Carbonneau et al. 2018) is a variant of a supervised machine learning method. In traditional supervised learning, an instance refers to a single data point. Each instance has one corresponding label. Whereas in MIL, labels are assigned to bags which are collections of instances, rather than to the individual instances within these bags. MIL is designed to leverage the detailed information in a bag to predict the label of the entire bag, even though individual instances inside the bag are not explicitly labeled. When predicting county-level corn yield by treating a county as an aggregation of all individual pixels, each county can be viewed as a bag and its comprising pixels can be treated as instances. To fully utilize the connection of instances in a bag, Cluster-MIL (Wagstaff, Lane, and Roper 2008) assumes that individual instances come from a group of underlying clusters, with a bag’s label being a function of one relevant cluster. The model learns the internal structure of bags that contain instances from various unknown distributions. Prime-MIL (Ray and Page 2001) supposes there exists a primary instance within each bag, then aims to select the most important instance for prediction. In contrast, Instance-MIL (Ins-MIL) (Ray and Craven 2005) employs each instance for regression directly, without taking the bag structure into account. There are two reasons that motivate us to investigate using MIL for county-level corn yield prediction modeling. First, MIL can make use of the detailed information in the pixels within a county for yield prediction, potentially utilizing underlying patterns within a county to assist the learning process. Second, in our scenario, only pixel-level RS imagery is available, but pixel-level yield records are lacking, making end-to-end pixel-level learning and analysis difficult. MIL can train models and make predictions even without instance labels, which avoids the end-to-end pixel-level learning.
When employing MIL to better utilize detailed information, we study satellite imagery at the pixel level and encounter a challenge commonly referred to as the “mixed pixel” problem (Xu et al. 2005), which arises from the inconsistent resolution between crop masks and various feature datasets. We use the MODIS dataset (Justice et al. 1998; Schaaf and Wang 2015) as a primary example to illustrate this problem. Fig. 1 (a) shows a $500\mathrm{m}\times 500\mathrm{m}$ MODIS pixel, which contains a patch of land with corn, bare soil, and weeds. Fig. 1 (b) displays the same MODIS pixel masked by the $30\mathrm{m}\times 30\mathrm{m}$ CDL crop mask (USDA-NASS 2017), where the green portion represents land planted with corn, and the blue portion indicates unrelated land types. This large MODIS pixel is inherently impure because it encompasses multiple land types. In our research, the primary focus is on the $30\mathrm{m}\times 30\mathrm{m}$ CDL cornfield pixels segmented by the CDL mask. When we study such a CDL cornfield pixel in the green area of Fig. 1 (b), its feature information, such as the Vegetation Indices (VIs), is derived from the encompassing large MODIS pixel, which contains a mixture of different land types. Therefore, even if this small CDL pixel is entirely composed of cornfield, its information is still mixed. This introduces noise into the models we build, affecting their performance. As a result, we refer to this small CDL cornfield pixel containing mixed information as a “mixed pixel”. In subsequent sections, any reference to a mixed pixel specifically indicates the small CDL cornfield pixels under investigation. It is worth noting that merely increasing the resolution of the crop mask does not solve the mixed pixel problem, given the inherent resolution limitations of satellite imagery. In addition, although we use the MODIS dataset as an illustration, the mixed pixel problem is not unique to satellite imagery; any feature with a resolution lower than that of the CDL mask can present this problem. To effectively address this issue, we propose using an attention mechanism (Vaswani et al. 2017) to rationally utilize the information within these mixed pixels.
Refer to caption(a)
Refer to caption(b)
Figure 1: The map from Google Earth of (a) One $500\mathrm{m}\times 500\mathrm{m}$ MODIS pixel and (b) this $500\mathrm{m}\times 500\mathrm{m}$ MODIS pixel masked by $30\mathrm{m}\times 30\mathrm{m}$ resolution CDL mask.
To address the “mixed pixel” problem, a straightforward approach is to calculate the mixing level of these mixed pixels and then directly use this mixing level to weight each pixel. However, this method poses certain challenges and limitations. When employing various datasets with different spatial resolutions, the choice of scale for calculating mixing level needs to be considered. It’s difficult to manually integrate the mixing levels from different datasets. Additionally, we also discuss the disadvantages of this method in detail in the discussion (Sec. 5). Therefore, we use an attention mechanism to automatically calculate the weights of mixed pixels. The attention mechanism (Vaswani et al. 2017; Ilse, Tomczak, and Welling 2018) is a powerful tool in deep learning. It enables models to focus on specific aspects of complex input data. Essentially, it can be used to assign different weights or “attention scores” to different parts of the input. This weighting mechanism is analogous to the way human beings pay attention to certain details while ignoring others (Khan et al. 2022). To address the “mixed pixel” problem, we propose to use this attention mechanism to allocate higher weights to pixels predominantly composed of cropland and lower weights to those with more irrelevant land types, thereby improving the model’s accuracy and interpretability. This method effectively utilizes the information from each pixel while mitigating the negative impact of noisy information in mixed pixels.
Many previous studies apply the combination of the attention mechanism and MIL. For example, Ilse, Tomczak, and Welling (2018) introduces a MIL method that is completely driven by neural networks, which includes trainable MIL pooling with the gated attention mechanism. Rymarczyk et al. (2021) incorporates self-attention into the process of MIL Pooling, merging the multi-level dependencies among different image areas with a trainable weighted average pooling mechanism. Hu et al. (2023) proposes a new MIL technique for medical image analysis named triple-kernel gated attention-based MIL with contrastive learning instead of pre-training. While the above studies have explored the combination of attention and MIL extensively, this study aims to effectively utilize detailed information through MIL and subsequently resolve the mixed pixel problem with an attention mechanism, which are two progressive objectives. Furthermore, in studies on the mixed pixel problem, no one has directly applied attention to this issue (Xu et al. 2005; Hsieh, Lee, and Chen 2001). To the best of our knowledge, this is the first application of the attention mechanism to address the mixed pixel problem in county-level corn yield prediction.
This paper presents a new method to enhance county-level corn yield prediction. Previous work mostly aggregated features to the county level, thereby ignoring pixel-level information. Our method utilizes the MIL framework to extract detailed information from individual pixels within each county, thereby enhancing performance. Furthermore, to solve the “mixed pixel” problem, our method uses the attention mechanism, which automatically assigns weights to each pixel based on its contribution to the final prediction. In the experimental validation, we not only achieved good results using our method on data from the U.S. Corn Belt but also verified the reliability of the results using some visualization techniques. To demonstrate the superiority of our method, we compared it with traditional machine learning methods and several promising deep learning models for county-level corn yield prediction. Currently, apart from the classic MLP (Haykin 1998), CNN (LeCun, Bengio et al. 1995), and LSTM (Hochreiter and Schmidhuber 1997)-based methods, the most popular deep learning model in remote sensing is the Transformer-based (Vaswani et al. 2017) model (Liu et al. 2022b; Lin et al. 2023; Bi et al. 2023). By comparing with these baseline models, we confirm that our method can effectively address the problems above and achieves better results.
In summary, the contributions of this paper are:
  • We fully utilize the detailed pixel-level information in RS data to improve prediction accuracy.
  • We use an attention mechanism to effectively solve the “mixed pixel” problem.
  • Our model achieves promising results on U.S. Corn Belt data, demonstrating its superiority over traditional machine learning and other deep learning models.
This paper is organized as follows. Sec. 1 introduces the background on corn yield prediction, the challenge, and the objective of this paper. Sec. 2 presents the study area, input data, experiment design, and the evaluation metrics. Sec. 3 describes the overview of the proposed attention-weighted MIL method including problem definition, attention mechanism, and MIL sampling method. Sec. 4 shows the results of the experiment, discusses some abnormal results in extreme weather, and gives some detailed analysis from both spatial and temporal perspectives. Sec. 5 delves into attention analysis and correlation of attention and agricultural features. Finally, Sec. 6 draws conclusions of this research.
2 Data acquisition
In our experiment, features are drawn from a range of sources (Table 1), including satellite imagery, weather, and soil properties, etc. We also incorporate the prediction year as a feature to enhance the model’s understanding of temporal patterns, and use the historical average yield to serve as a baseline to improve the model’s robustness.
Category Variables Unit Related Properties Spatial Resolution Source Latency
[…]
Green Chlorophyll Index (GCI)
Normalized Difference Water Index (NDWI)
[…]
Nighttime Land Surface Temperature (LSTnight)
[…]
Max Temperature (Tmax)
Minimum Temperature (Tmin)
Total Precipitation (PPT) $\mathrm{mm}$ Water stress
maximum vapor pressure deficit (VPDmax) $\mathrm{hPa}$
minimum vapor pressure deficit (VPDmin)
[…]
Soil Organic Matter (SOM) $\mathrm{kg}$/$\mathrm{m}^{2}$ Soil nutrient uptake
Cation Exchange Capacity (CEC) $\mathrm{cmol}$/$\mathrm{kg}$
Others Historical average yield t/ha N/A County-level USDA NASS
Prediction year N/A N/A
Table 1: This table summarizes all the data information used in our experiments.
2.1 Study area
The study area covers twelve states in the U.S. Corn Belt: North Dakota, South Dakota, Minnesota, Wisconsin, Iowa, Illinois, Indiana, Ohio, Missouri, Kansas, Nebraska, and Michigan. We calculate the average corn yield for all states from 2008 to 2022, as shown in Fig. 2. Darker colors represent higher yields. The corn yield in the selected states accounts for most of the US total corn yield (USDA 2020b), which is sufficient to support our predictions of corn yield for the United States. This experimental setting follows several previous works (Ma and Zhang 2022; Ma et al. 2021b, 2019).
Figure 2: The 5-year average yield map of each county.
2.2 Satellite data
[…]
$GCI=\frac{NIR}{Green}-1$ (1)
Where $NIR$ represents the surface reflectance in the Near Infrared band, and $Green$ represents the surface reflectance in the Green band.
EVI is an enhanced vegetation index, which optimizes the vegetation information by minimizing soil and atmosphere influences. It is particularly effective in areas with dense vegetation and provides more details about biomass and canopy structural variations:
$EVI=2.5\times\frac{NIR-Red}{NIR+6\times Red-7.5\times Blue+1}$ (2)
Where $Red$ and $Blue$ represent the surface reflectances of the Red band and Blue band respectively.
NDWI, the normalized difference water index, is used to monitor changes in the water content of leaves, helping to identify vegetation stress due to drought or disease. This index is particularly useful in monitoring crop conditions where precise water management is crucial:
$NDWI=\frac{NIR-SWIR}{NIR+SWIR}$ (3)
Where $SWIR$ represents the short-wave infrared band.
2.3 Weather data
The Parameter elevation Regressions on Independent Slopes Model (PRISM) dataset (Daly, Smith, and Olson 2015; Daly et al. 2008) is a highly sophisticated climate data product that offers detailed information on various climatic elements, including precipitation, temperature, and humidity, across the United States. The weather observations used for the experiment include daily mean air temperature (Tmean), maximum air temperature (Tmax), minimum air temperature (Tmin), maximum vapor pressure deficit (VPDmax), minimum vapor pressure deficit (VPDmin), and total precipitation (PPT) from the PRISM dataset with 4 $\mathrm{km}$ spatial resolution. Tmean represents the average air temperature over a 24-hour period. Tmax and Tmin represent the highest and lowest air temperatures recorded in a day, respectively. VPDmax and VPDmin reflect the maximum and minimum differences respectively, between the amount of moisture in the air and the amount it can hold when it is saturated. PPT indicates the total amount of rain, snow, or other precipitation falling in a given time period. Moreover, daytime and nighttime land surface Temperature (LSTday and LSTnight) are extracted from the MODIS MYD11A2 product (Park, Feddema, and Egbert 2005) with 1 $\mathrm{km}$ spatial resolution. LSTday and LSTnight represent the temperature of the Earth’s surface during the daylight and night hours. These meteorological features are crucial in understanding weather patterns and can have significant impacts on agricultural and environmental conditions.
2.4 Soil data
[…]
2.5 Cornfield mask
[…]
2.6 Data preprocessing
On the Google Earth Engine (GEE) platform (Gorelick et al. 2017), we spatially aggregate all the RS and weather features (GCI, EVI, NDWI, LSTday, LSTnight, Tmean, Tmax, Tmin, PPT, VPDmin, VPDmax, AWC, SOM, CEC). These features are then grouped temporally into 16-day intervals from mid-May to early October to cover the corn-growing period. The dataset covers the years 2008 to 2022 and includes county-level historical yield records obtained from the USDA NASS (USDA 2020b). Furthermore, our model integrates the year data to capture specific time-related features and calculates the 5-year historical average yield to provide a baseline for model robustness. The feature length of each input data point is 159. After processing, the dataset includes approximately 600 to 800 counties per year, with each county represented by 100 pixels. The test set is from 2018 to 2022, while the training set ranges from 2008 to the year preceding the testing year. We set aside 20% of the training data for model validation.
3 Methodology
In Fig. 3, our method is structured as follows: we begin by applying the CDL mask to the MODIS dataset. This is followed by selecting some pixels within a county for use in training the model and making predictions. Next, we employ MIL to establish a many-to-one mapping. Finally, we implement an attention mechanism to assign weights to these pixels, expecting to assign higher weights to pure pixels and lower weights to mixed pixels.
Figure 3: The pipeline of the proposed method.
In this section, we will 1) formally define the county-level corn yield prediction (Sec. 3.1); 2) describe the MIL method and how to apply it to county-level corn yield prediction with individual pixels’ information input (Sec. 3.2); 3) introduce the attention mechanism in detail and explain its application in our setting (Sec. 3.3).
3.1 Problem definition
[…]
To formalize the county level crop yield prediction problem, let $n$ be the number of counties in the given study area. Let $I_{i}$ be the satellite imagery and $y_{i}$ be the yield for each county $i$, $i=1,...,n$. Each image is on the space $\mathbb{R}^{W\times H\times C}$, where $W$ and $H$ are the width and height of the image, and $C$ denotes the channels. We calculate VIs from the satellite imagery, along with other agricultural and environmental features (weather, soil properties, etc.), to form our county-level input feature $F_{i}$. From a machine learning perspective, our training objective is to develop a model $f$, $f(F_{i})=y_{i}$.
3.2 Multiple instance learning
Conventional county-level corn yield prediction methods usually aggregate all the pixels in one county into a single value (Ma, Yang, and Zhang 2023; Ma et al. 2023, 2024), which may compromise the integrity of county data representation, potentially losing some detailed information. In response to these limitations, we investigate pixel-level feature processing and apply MIL to the task. In real-world applications, there are often scenarios where a single label represents a group of data points. In MIL, this setting is described as a bag containing many instances, from which we need to predict the label. Thus, unlike the traditional one-to-one mapping in machine learning, MIL involves a many-to-one mapping. Additionally, MIL usually assumes that there is some connection between instances within a bag or that the bag has some kind of special structure. e.g., to fully utilize the connection of instances in a bag, Cluster-MIL (Wagstaff, Lane, and Roper 2008) assumes that individual instances come from a group of underlying clusters and employs the MIL model to learn the internal structure of bags that contain instances from various unknown distributions.
In our work, we follow the approach of (Ilse, Tomczak, and Welling 2018), using an attention module to process each county and learning the connections between pixels within the county to improve the model performance. One key difference is that their work focuses on classification, whereas ours focuses on regression. We will first provide a detailed description of how attention is used to perform the MIL classification task.
Problem formulation for MIL in classification task
In traditional binary supervised learning, the goal is to find a model that predicts the value of a target variable, $y\in\{0,1\}$, for a given instance, $\mathbf{x}\in\mathbb{R}^{D}$. However, in the MIL problem there is a bag of instances, $X=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\}$. Each bag has a single binary label $Y$ associated with it. Additionally, It is assumed that individual labels exist for the instances within a bag, i.e., $y_{1},\ldots,y_{K}$ and $y_{k}\in\{0,1\}$, for $k=1,\ldots,K$, but these labels are inaccessible and remain unknown during training.
The assumptions of the MIL problem can be expressed as follows:
$Y=\begin{cases}0,&\text{if}\ \sum_{k}y_{k}=0,\\ 1,&\text{otherwise}.\end{cases}$ (4)
or it can be re-formulated as:
$Y=\max_{k}\{y_{k}\}.$ (5)
These assumptions imply that an MIL model must be permutation-invariant (Zaheer et al. 2017; Qi et al. 2017). The key component in constructing an MIL model is to combine the instances with a permutation-invariant function $\sigma$, which is also referred to as MIL pooling.
Ilse, Tomczak, and Welling (2018) proposed using an attention module (Vaswani et al. 2017) to weight the instances using neural network. We will explain the attention mechanism in detail later (Sec. 3.3). The weighted average function meets the requirements for being permutation-invariant. Given a bag $X$ of $K$ instances, $X=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\}$, the following MIL pooling ($\sigma$) is proposed:
$\mathbf{z}=\sum_{k=1}^{K}a_{k}\mathbf{x}_{k},$ (6)
where
$a_{k}=\frac{\exp\left(\mathbf{w}^{\top}\tanh\left(\mathbf{V}\mathbf{x}_{k}^{\top}\right)\right)}{\sum_{j=1}^{K}\exp\left(\mathbf{w}^{\top}\tanh\left(\mathbf{V}\mathbf{x}_{j}^{\top}\right)\right)},$ (7)
where $\mathbf{w}\in\mathbb{R}^{L\times 1}$ and $\mathbf{V}\in\mathbb{R}^{L\times M}$ are parameters. Moreover, the hyperbolic tangent $\tanh(\cdot)$ element-wise non-linearity is used for both positive and negative values, ensuring proper gradient flow. This construction enables the discovery of connections among instances. Afterwards, when the permutation-invariant function $z$ has been constructed , any classification module can be applied to predict the class of the bag. Ideally, when the label $Y=1$, high attention weights should be assigned to instances likely to have the label $y_{k}=1$.
Problem formulation for MIL in regression task For the county-level corn yield prediction task, we have the yield (label) of each county (bag), and each county is composed of a number of pixels (instances). By utilizing the detailed information of each county, MIL can improve prediction accuracy. Suppose there are $m_{i}$ pixels in each county $i$, $i=1,...,n$. To differentiate from the county-level feature $F$, we use $P$ to represent pixel-level features. For each pixel $j$, $j=1,...,m_{i}$, we extract its feature $P^{j}_{i}$. There exists a model $f$ with many-to-one mapping, namely $f(P^{1}_{i},P^{2}_{i},...\ldots,P^{m_{i}}_{i})=y_{i}$. We hope our model can automatically learn the structure of a county as well as the connections between different pixels within it, thereby improving corn yield predictions.
A key step in MIL is selecting representative instances within each bag, since training on all instances in a bag is computationally expensive, especially for large bags. Additionally, without clear indications of which instances are more influential for a bag’s label, treating all instances with equal importance could mislead the model during the learning process. In our county-level corn yield prediction task, we utilize a unique sampling method, drawing inspiration from the mixed pixel problem, to select the most representative pixels within each county. As illustrated in Fig. 1, the VIs of a CDL pixel are derived from the information of the encompassing MODIS pixel; all the CDL pixels within one MODIS pixel share the same VI. We believe that within a large MODIS pixel, we only need to select one CDL pixel to represent it; the rest are redundant. Therefore, in the MIL sampling process, we randomly select one CDL pixel from each MODIS pixel, removing the remaining pixels with duplicate VIs.
To simplify the learning process, we need to sample an appropriate number of pixels in each county. On the one hand, selecting fewer pixels might not capture the complete information in each county. On the other hand, over-sampling may result in an insufficient number of qualified counties, making it impossible to obtain sufficient data for training. Here, we randomly select 100 pixels from each county $i$, $i=1,...,n$, denoted as $\hat{P_{i}}=(P^{1}_{i},P^{2}_{i},...,P^{100}_{i})$. Our experiments in Section 4.4 have shown that these selected pixels will adequately capture the county’s information, as will be demonstrated in the experiment (Sec. 4.4). For simplicity of notation, we omit the subscript $i$ of $\hat{P_{i}}$ here, so it becomes $\hat{P}$. We aim to train a model $f$, $f(\hat{P})=y_{i}$, that can predict the yield accurately.
3.3 Attention mechanism implementation
In previous sections, we discuss the mixed pixel problem, where individual CDL pixels contain a mixture of different land types information. This issue is particularly relevant when dealing with datasets of varying spatial resolutions, making it difficult to manually integrate and weight these mixed pixels effectively. To overcome these challenges, we employ an attention mechanism. The attention mechanism automatically assigns higher weights to pixels whose information is predominantly composed of cornfield and lower weights to those with more irrelevant land types.
The attention mechanism (Vaswani et al. 2017) in machine learning assigns varying importance to different input elements, enabling the model to focus on inputs that are more relevant to the output (Khan et al. 2022). This helps models perform better in complex tasks by learning context-specific representations. In our task, we not only use attention to implement MIL, but also to assign appropriate weights to the pixels in each county, hoping that attention can automatically determine the significance of each pixel. Ideally, it assigns higher weights for pure pixels and lower weights for mixed pixels. Suppose there are $100$ pixels in county $i$, $i=1,...,n$. For the feature $P_{i}^{j}$ of pixel $j$, $j=1,...,100$, our goal is to learn a weight $W_{i}^{j}$. We are aiming for such a model $f$, $f({W^{1}_{i}}^{\top}P^{1}_{i},{W^{2}_{i}}^{\top}P^{2}_{i},...,{W^{100}_{i}}^{\top}P^{100}_{i})=y_{i}$. In this way, we can reduce as much the noisy information in mixed pixels as possible.
The neural network learns patterns from input data through a process called forward propagation by some layers and performs error adjustment with a backpropagation process (Goodfellow, Bengio, and Courville 2016). In the attention module, we use $100$ pixels from each county $i$ and extract their features $P^{j}_{i}$ as the input vector $\hat{P_{i}}=(P^{1}_{i},P^{2}_{i},...,P^{100}_{i})$. For simplicity of notation, we omit the subscript $i$ of $\hat{P_{i}}$ here, so it becomes $\hat{P}$. Then it is passed through three separate modules $f_{Q}$, $f_{K}$, and $f_{V}$ to obtain the respective vectors: query $Q$, key $K$, and value $V$. The three separate modules are fully connected layers, which are a fundamental building block in artificial neural networks. In these layers, each neuron is connected to every neuron in the previous layer. Each connection has an associated weight, which determines the strength and direction (positive or negative) of the influence of the input on the output. Additionally, each output neuron has a bias term that is added to the weighted sum of inputs before applying the activation function. The output of a fully connected layer is computed by taking the dot product of the input vector and the weight matrix, adding the bias vector, and then applying an activation function.
$Q=f_{Q}(\hat{P}),K=f_{K}(\hat{P}),V=f_{V}(\hat{P})$ (8)
Then we calculate the attention score on a batch of queries $Q$, keys $K$, and values $V$ concurrently, the output matrix is computed as follows:
$Attention(Q,K,V)=softmax(\frac{QK^{\top}}{\sqrt{d_{K}}})V$ (9)
Where $d_{K}$ is the dimension of $K$.
And The softmax function (Goodfellow, Bengio, and Courville 2016) is defined as:
$\text{softmax}(\mathbf{z})_{i}=\frac{e^{z_{i}}}{\sum_{j=1}^{K}e^{z_{j}}}$ (10)
Where $\mathbf{z}=(z_{1},\dots,z_{K})$ is the input vector for $i=1,\dots,K$. $e^{z_{i}}$ is the exponential of the $i$-th element of the input vector $\mathbf{z}$.
The similarity between the query $Q$ and the key $K$ is captured through their dot product. The higher the dot product, the more similar the query $Q$ is to the key $K$. The dot product result is then passed through a softmax function to obtain the weights, which range between 0 and 1 and sum to 1. The softmax function ensures that the larger dot products get higher weights. Finally, these softmax weights are used to compute a weighted sum of the values $V$. Each value is multiplied by its corresponding weight, and these products are then summed. This final result is the output of the attention mechanism, which represents the context relevant to the query $Q$ given key-value pairs.
By combining the above equations, we denote the function of our attention module as $Att(\hat{P})$. Then our model $f$ becomes $f(Att(\hat{P}))=y_{i}$.
Using attention to weight the pixels in each county allows us to achieve MIL, enabling the learning of each county’s detailed information. This method also helps in determining the completeness of information from different pixels, effectively filtering out noisy data in mixed pixels.
4 Experiment
4.1 Experimental setup
In order to evaluate the performance of the proposed model, three traditional machine learning models, including linear regression (LR) (Bishop and Nasrabadi 2006), ridge regression (RR) (Hoerl and Kennard 1970), and random forest (RF) (Breiman 2001), are selected as baseline models for comparison. In particular, LR is a statistical method for predicting an outcome based on the linear relationship with one or more independent variables. RR is a variant of LR that adds a penalty term to prevent overfitting in high-dimensional datasets. RF is an ensemble machine learning algorithm that aggregates many decision trees to achieve more accurate predictions. These models are implemented using scikit-learn (Pedregosa et al. 2011) with default parameters. To highlight the role of MIL, Ins-MIL (Ray and Craven 2005) is also included as a comparison, which employs each instance for regression directly, without taking the bag structure into account. Additionally, based on our review of the application of deep learning models in remote sensing (Sec. 1), we select the current mainstream models—CNN (LeCun, Bengio et al. 1995), LSTM (Hochreiter and Schmidhuber 1997), ResNet He et al. (2016), and Transformer (Vaswani et al. 2017). Specifically, for the CNN model, since we are not directly inputting 2D remote sensing images but rather a temporal vector, we replace the traditional 2D convolution with 1D convolution. For LSTM, ResNet, and Transformer, we use PyTorch’s built-in model (Paszke et al. 2019) and adjust the input and output to fit our data format. To use the PyTorch built-in ResNet18 model with 1D data, we adapt the 1D input to match the expected 2D input format, which is originally designed for images. This involves reshaping the input data, modifying the first convolutional layer, and replacing it to accept a single input channel. For all the deep learning models we use Adam (Kingma and Ba 2014) as the optimizer with a batch size of $100$. Initially, the learning rate is set to $0.1$ and is reduced by a factor of 10 whenever the error reaches a plateau. All experimental results are obtained after averaging five repetitions. For each experiment, we randomly select different random seeds to ensure the results are fair and reproducible.
4.1.1 Performance Metrics
The experiment results are measured by standard error metrics $RMSE$ and $R^{2}$ (Albergel et al. 2013):
$RMSE=\sqrt{\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{n}}$ (11)
$R^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})_{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}$ (12)
[…]
4.2 Model Performance Evaluation
The proposed model, MIL with attention mechanism (Att-MIL), is compared with other methods in Table 2. The results show that our approach outperforms all other methods across all years and metrics. The superior performance of our approach can be attributed to its effective extraction and incorporation of inter-pixel correlations within each county, which other methods overlook. The second-best method, Ins-MIL, effectively utilizes the MIL concept and recognizes that each county consists of multiple pixels. This understanding provides a more comprehensive view than merely considering each county as a single entity and results in better performance compared to traditional regression-based methods. For the traditional machine learning models, RF, ranking third, exhibits relatively good performance due to its robustness to noise and ability to model non-linear relationships. However, it does not inherently model detailed county information, thereby performing worse than the MIL-based methods. LR and RR are based on the assumptions of a linear relationship, which are not entirely accurate in the context of county-level corn yield prediction. These two methods exhibit the weakest performance in our experiments.
For the deep learning-based models, ResNet18 performs the worst. This is likely due to the significant difference between the pre-trained dataset and our dataset, as well as the fact that ResNet is not designed for 1D features. The CNN model performs slightly better. Although we modify the 2D convolution to 1D to adapt to our input data, the focus of CNN on local correlations limits its effectiveness on our dataset. The Transformer model comes next, as its large architecture requires substantial data to avoid overfitting, and our data volume is clearly insufficient. The LSTM model performs the best and is closest to the performance of our method. LSTM model excels in capturing long-term temporal dependencies and sequential patterns, which are crucial for yield prediction, as seasonal and temporal variations significantly impact results. The above comparisons underscore the benefits of integrating the attention mechanism and MIL in tackling the complexity of county-level corn yield prediction task.
Year Method Att-MIL Ins-MIL LR Ridge RF
$RMSE$ $R^{2}$ $RMSE$ $R^{2}$ $RMSE$ $R^{2}$ $RMSE$ $R^{2}$ $RMSE$ $R^{2}$
2018 1.00 0.67 1.42 0.50 1.49 0.45 1.49 0.45 1.41 0.50
2019 0.99 0.58 1.18 0.51 1.46 0.26 1.45 0.27 1.60 0.10
2020 1.00 0.48 1.35 0.20 1.53 -0.02 1.81 -0.42 1.41 0.13
2021 0.99 0.64 1.16 0.60 1.78 0.05 2.04 -0.24 1.08 0.64
2022 0.83 0.84 1.07 0.73 1.35 0.65 1.18 0.73 1.19 0.73
Year Method CNN LSTM Resnet18 Transformer
$RMSE$ $R^{2}$ $RMSE$ $R^{2}$ $RMSE$ $R^{2}$ $RMSE$ $R^{2}$
2018 1.48 0.42 1.03 0.62 1.67 0.42 1.12 0.55
2019 1.34 0.44 1.01 0.52 1.53 0.32 1.15 0.47
2020 1.33 0.21 1.02 0.42 1.73 0.27 1.07 0.41
2021 1.41 0.47 1.06 0.57 1.67 0.35 1.14 0.51
2022 1.42 0.62 0.91 0.79 1.32 0.45 0.97 0.77
Table 2: Model evaluation results in 2018-2022.
To illustrate the relationship between predicted and reported yields, we use scatter plots to provide a clear visual representation of their alignment (Fig. 4). Each point in the figures corresponds to a specific county. From these scatter plots, we observe that predictions made by Att-MIL are tightly gathered around the line of perfect fit. The scatter plot of Ins-MIL, although less gathered than that of our method, still demonstrates a significant degree of accuracy. RF does not capture detailed information within counties; hence, its results on the scatter plot are more scattered compared to the MIL-based methods. In contrast, LR and RR display the greatest spread of data points. This indicates that our approach provides more consistent and accurate results and outperforms the other methods.
Figure 4: Scatter plots of observed yields vs. predicted yields of (a) Att-MIL, (b) Ins-MIL, (c) LR, (d) Ridge, (e) RF in 5 testing years: (1) 2018; (2) 2019; (3) 2020; (4) 2021; (5) 2022. The colorbar represents the density of points at each location.
Next, to visualize the magnitude of prediction errors spatially across different regions, we present a choropleth map that illustrates the absolute error of each method across various states (Fig. 5). The absolute error map provides a geographic representation of the prediction performance of each method at the county level. On the map, each colored block represents a county, with color intensity indicating the prediction error magnitude, deeper colors signify higher errors. Upon examination, it becomes clear that our method presents the lightest color across most counties. This suggests that our approach achieves the lowest absolute error overall, indicating superior prediction performance.
Figure 5: Absolute error maps of (a) Att-MIL, (b) Ins-MIL, (c) LR, (d) Ridge, (e) RF in 5 testing years: (1) 2018; (2) 2019; (3) 2020; (4) 2021; (5) 2022.
[…]
Figure 6: The error maps in some abnormal years: (a) Ins-MIL in 2020, (b) LR in 2020, (c) LR in 2021, (d) Ridge in 2020, (e) Ridge in 2021. A negative value indicates overestimation, while a positive value indicates underestimation.
4.3 In-season yield prediction
To verify the capability of our model in providing good prediction throughout various growth stages of corn, we conduct a comparative experiment of all models for in-season county-level corn yield prediction. The timeframe is divided from mid-May to early October into 10 intervals, predicting once every 16 days (Fig. 7). Predicting corn yield is challenging in the initial stages of the growing season because the correlation between satellite imagery features and yield is very weak (Johnson 2014b). This experiment is based on the data from 2022, with the training data spanning from 2008 to 2021. It is clear that our method incorporating attention and MIL achieves the best $RMSE$ and $R^{2}$ values, indicating it provides the best fit to the data across all dates. In the mid-term stage of corn growth, the performance of the four methods other than Att-MIL is roughly similar. Ins-MIL performs second-best for most of the time, while RF is somewhat inferior. LR and RR perform the worst in the early and late stages. This superior performance of our model is due to its ability to manage uncertainty and noise during the early stages of corn growth.
Figure 7: In in-season county-level corn yield prediction in 2022: (a) $RMSE$ and (b) $R^{2}$.
4.4 The number of instances in a bag
In our experiment, a crucial parameter is the number of pixels selected from each county. Selecting an excessive number of pixels in each county may lead to a reduced number of qualified counties. Conversely, if the number of selected pixels is too small, they may not adequately represent the complete information of a county, negatively impacting the model’s training and prediction. This naturally introduces a trade-off where more pixels in a single county lead to fewer counties being selected, and vice versa. The objective of this section is twofold: (1) identifying an optimal parameter range that allows the model to reduce the size of the dataset without sacrificing accuracy; and (2) studying the model’s robustness under various parameter conditions. Therefore, we conduct a thorough experiment by varying the number of pixels per county between 2 and 1500 (Fig. 8). This experiment is based on the testing years from 2018 to 2022, while the training year ranges from 2008 to the year preceding the testing year.
Figure 8: The results of our model in 2022 under different numbers of selected pixels in one county: (a) $RMSE$ and (b) $R^{2}$.
As shown in Fig. 8, our proposed Att-MIL model consistently demonstrates superior performance, while the Ins-MIL and RF models exhibit significantly poorer performance. Meanwhile, LR and RR perform substantially worse as the number of pixels per county increases (RMSE increases and R2 decreases). This might be attributed to these models’ limited ability to learn from larger sample datasets . Notably, even as the pixel number of a county decreases to around 100, our model’s predictive performance remains robust. In addition, the results suggest a suitable pixel count ranges from 100 to 1000. Within this range, the predictions of our Att-MIL model exhibit minimal fluctuation, with the optimal outcome achieved around 100 pixels. These findings highlight the effectiveness and robustness of the proposed Att-MIL model, particularly in scenarios with limited data. This implies that our model is well-suited for handling datasets of varying sizes, demonstrating a significant advantage in real-world county-level corn yield prediction.
5 Discussion
This section provides a detailed analysis of the attention mechanism. In Sec. 5.1, we prepare the necessary data and perform data preprocessing for the analysis experiment. Next, in Sec. 5.2, we pose two important questions regarding the attention mechanism. In Sec. 5.3, we delve into a deep analysis of attention by answering the two proposed questions. Then, in Sec. 5.4, we select a few counties as examples, magnifying their satellite images to demonstrate the role of attention. Furthermore, we summarize the advantages of attention in Sec. 5.5. Finally, we discuss and explain the limitations of our method (Sec. 5.6).
5.1 Pixel mixing level calculation and preprocessing
To explore the role of attention, we first need to quantify the mixed pixels based on their definition, which is the proportion of the cornfield within a pixel of the feature dataset. Taking the MODIS dataset as an example, suppose one MODIS pixel contains $n$ CDL pixels. We represent each MODIS pixel as a binary list $S$ of length $n$, where each element $s_{i}$ represents a CDL pixel, $s_{i}\in\{0,1\}$ for $i=1,2,...,n$. Then, using GEE, we calculate the corn ratio, defined as the number of CDL pixels that represent corn within a MODIS pixel, to serve as the ground truth for our experiment:
$corn\>ratio=\frac{\Sigma_{i=1}^{n}s_{i}}{n}$ (13)
For the following analysis, the initial step of preprocessing is “bag normalization” by normalizing the corn ratio within each bag so that their sum equals one. Suppose we select $m$ pixels in each county $i$, $i=1,...,n$. In county $i$, we have a corn ratio $C_{i}^{j}$ for each pixel $j$, $j=1,...,m$. Then, the implementation of bag normalization is as follows:
${C_{i}^{j}}^{\prime}=\frac{C_{i}^{j}}{\Sigma_{k=1}^{m}C_{i}^{k}}$ (14)
Where ${C_{i}^{j}}^{\prime}$ is bag-normalized corn ratio.
The purpose of bag normalization is to make attention and the corn ratio comparable. We first normalize the corn ratio values within each bag, transforming them into relative values. This ensures the fairness of subsequent experiments.
5.2 Two questions towards attention mechanism
In the above experiments, we demonstrate that Att-MIL can indeed enhance the predictive ability of the model. To ascertain the specific role that attention plays within the model, further in-depth analysis is necessary. We need to answer two questions: (1) Which part of the input contributes the most? and (2) Do the weights learned by the attention mechanism truly reflect the degree of mixing in each pixel? In this section, we will demonstrate the specific function of attention by analyzing the relationship between attention, features, and corn ratio. This experiment is based on the data from 2022, with the training data spanning from 2008 to 2021.
5.3 Attention’s correlation with features and corn ratio
5.3.1 Feature importance
To answer the first question, examining the importance of features is essential because we can verify whether the model derives attention based on relatively important features (Khan et al. 2022). To investigate feature importance, we employ the feature ablation strategy (Litkowski 2016). For each feature, we retrain our model after excluding that feature to test its impact on the model’s performance (Table 3). In this way, we can infer that if there’s a notable decline in performance when a feature is removed, it indicates the feature’s crucial significance.
Metrics Feature All SOM Year CEC PPT VPDmin VPDmax Tmin LSTnight
$RMSE$ 0.830 0.848 0.865 0.875 0.889 0.901 0.928 0.934 0.954
$R^{2}$ 0.840 0.832 0.829 0.829 0.818 0.814 0.814 0.809 0.799
Metrics Feature AWC Tmean NDWI EVI LSTday GCI Tmax Historical
$RMSE$ 0.975 0.992 1.003 1.024 1.051 1.059 1.062 1.130
$R^{2}$ 0.790 0.789 0.784 0.783 0.780 0.777 0.774 0.759
Table 3: This table displays the results of our feature ablation experiments. All features are arranged in ascending order of importance, from left to right and from top to bottom.
Through this examination, we found that the descending order of feature importance is as follows: historical average yield, Tmax, GCI, LSTday, EVI, NDWI, Tmean, AWC, LSTnight, Tmin, VPDmax, VPDmin, PPT, CEC, year, and SOM (Fig. 9). The historical average yield is considered to be of 100% importance, and the least important feature SOM is treated as 0% importance. All the features are arranged based on their experimental metrics. The importance of the other features is calculated using the formula:
Figure 9: The normalized importance ranking of a total of 16 features.
$\text{Importance}=\left(\frac{\text{value}-\text{min}}{\text{max}-\text{min}}\right)\times 100\%$ (15)
where $value$ is the metric ($RMSE$ and $R^{2}$) of the feature being evaluated, $min$ is the metric of the least important feature (SOM), and $max$ is the metric of the most important feature (historical average yield). This formula scales the importance of each feature relative to the range between the minimum and maximum metric values.
The most important feature, historical average yield, suggests that past yield data is crucial in predicting future yields. This aligns with the idea that past performance can often provide a strong baseline for future prediction. Tmax and GCI rank high, implying the importance of temperature and vegetation health in corn yield. On the other hand, the relatively low importance of features like year and SOM indicates that these features, while having some predictive power, contribute less significantly compared to other powerful features. We rank the importance of features here to delve deeper into their relationship with attention in subsequent analyses (Sec. 5.3.2).
5.3.2 Attention’s correlation with input features
Since our model’s input consists of various features, we aim to understand how the model learns attention from these features using the Pearson correlation coefficient (Freedman, Pisani, and Purves 2007). The Pearson correlation coefficient quantifies the linear relationship between two continuous variables. We identify some of the most important features in the feature importance experiment (Sec. 5.3.1) and explore how the model learns attention by calculating the correlation coefficients between these features and attention. We perform the same pre-processing “bag normalization” as in Sec. 5.1 to convert it into relative value.
As displayed in Table 4, the results indicate that attention has a high correlation with several important features, such as VIs, LSTday, and Tmax. In the context of statistical analysis, a correlation coefficient of 0.5 is significant enough to confirm that there is a strong correlation between the two factors (Schober, Boer, and Schwarte 2018). Since we know that the function of attention is to learn the part of the input that has a higher correlation with the output, we can conclude that the attention mechanism assigns attention scores by learning those important features.
Variables NDWI GCI EVI LSTday Tmax Corn ratio
Correlation 0.50 0.53 0.48 0.45 0.52 0.68
Table 4: The Pearson correlation coefficient between attention with some of the most important features and corn ratio.
5.3.3 Attention’s correlation with corn ratio
To answer the second question, we explore whether the attention truly reflects the degree of mixing in mixed pixels, thereby assigning the correct weights to these pixels. Our goal is to empirically demonstrate the connection between the attention and the corn ratio. As displayed in Table 4, the results suggest a high correlation between the attention score and the corn ratio, indicating that our attention mechanism accurately predicts the degree of mixing in the mixed pixels.
5.4 Zoomed-in mixed and pure pixels analysis
Previous experiments validated the roles of attention, corn ratio, and features from a statistical perspective. Based on our experimental setup, 100 CDL pixels are randomly selected in each county to represent the features of that county. We delve into each county to visually examine their connections (Fig. 10). Four counties are selected from all the data as examples from top to bottom. From left to right, the figures represent the attention map (a), corn ratio (b), a zoomed-in view of pure pixels (c), and a zoomed-in view of mixed pixels (d). In the visualization graphs of attention and corn ratio, the color of each point represents the magnitude of the value. The deeper the color, the larger the value. In the zoomed-in view of pixels, each small red dot represents the location of a CDL pixel. Furthermore, we aim to determine whether a CDL pixel is mixed or pure based on the surrounding land types. To this end, we plot $500\mathrm{m}\times 500\mathrm{m}$ red squares around the CDL pixels’ locations to check whether the areas surrounding these CDL pixels contain cornfields or other types of land.
First, we examine the county-level attention visualization results from a broad perspective. By comparing the attention maps (Fig. 10 a1-a4) with the corn ratio maps (Fig. 10 b1-b4), we aim to demonstrate that our attention mechanism accurately reflects the proportion of cornfield pixels in each county. In Fig. 10 (a1) and (b1), the top-left corner shows a cluster of darker-colored pure pixels, while the top-right corner features lighter-colored mixed pixels, indicating a strong correspondence between the attention and corn ratio. In Fig. 10 (a2) and (b2), the left half of the pixels are darker, and the upper right section is lighter. Similarly, in Fig. 10 (a3) and (b3), the center pixels are darker, while the right side pixels are lighter. In Fig. 10 (a4) and (b4), the bottom left and top right sections of the pixels are darker, while the diagonal from the top left to the bottom right is lighter. These patterns clearly illustrate the correlation between our attention mechanism and the corn ratio at the county level.
Refer to caption(a1)
Refer to caption(b1)
Refer to caption(c1)
Refer to caption(d1)
Refer to caption(a2)
Refer to caption(b2)
Refer to caption(c2)
Refer to caption(d2)
Refer to caption(a3)
Refer to caption(b3)
Refer to caption(c3)
Refer to caption(d3)
Refer to caption(a4)
Refer to caption(b4)
Refer to caption(c4)
Refer to caption(d4)
Figure 10: Visual examination of (a) attention map, (b) corn ratio map, (c) zoomed-in pure pixels and (d) zoomed-in mixed pixels in four areas (1-4) from top to bottom.
Next, we zoom into the clusters of pure pixels (c1-c4) and mixed pixels (d1-d4) to validate the above analysis. It is observed that in the pure pixels, all the land is cornfield, while in the mixed pixels, a significant portion of the land is occupied by unrelated categories such as forests. This fully demonstrates that our attention captures the information of the mixed cropland.
5.5 Attention’s advantage over corn ratio
A simple solution to the mixed pixel problem is to directly calculate the corn ratio for each mixed pixel, and use it to weight each pixel before predicting corn yield. We conducted an experiment here to demonstrate that this method is not a preferable solution. We calculate the corn ratio using the CDL mask and the MODIS dataset, and then explicitly apply this corn ratio as a weight to the features of each pixel. Suppose we select $100$ pixels in each county $i$, $i=1,...,n$. We have the corn ratio $C_{i}^{j}$, feature vector $P_{i}^{j}$, and corn yield $y_{i}$ for each pixel $j$, $j=1,...,100$. Furthermore, the implementation of manual weighting is as follows:
${P_{i}^{j}}^{\prime}=C_{i}^{j}P_{i}^{j}$ (16)
Where ${P_{i}^{j}}^{\prime}$ is the manually-weighted feature vector.
Next, we use this manually weighted pixel dataset to predict corn yield $y_{i}$, namely $f({P^{1}_{i}}^{\prime},{P^{2}_{i}}^{\prime},...,{P^{100}_{i}}^{\prime})=y_{i}$. The model backbone is similar to that of our Att-MIL model. The only difference is the absence of the attention module; the manually-weighted method directly multiplies the corn ratio onto the pixel features instead. The experimental results are shown in Table 5, which compares the results of manual weighting and automatic weighting (Auto-weighting) methods. We find that the results of using the corn ratio as weights are not ideal. This validates our argument that using attention to automatically learn the weights can achieve better results.
Year Method Auto-weighting Manual weighting
$RMSE$ $R^{2}$ $RMSE$ $R^{2}$
2018 1.00 0.67 1.35 0.47
2019 0.99 0.58 1.27 0.55
2020 1.00 0.48 1.43 0.31
2021 0.99 0.64 1.24 0.57
2022 0.83 0.84 1.17 0.69
Table 5: Comparison between manually calculating the mixing level of mixed pixels and applying weights versus automatically weighting with attention.
In addition to the significantly better experimental results as shown in Table 5, the attention method also has other benefits over manual weighting. First, there are inherent errors in the calculation of the corn ratio. This is attributed to the datasets’ pixel sizes not being perfectly divisible and the boundaries of various pixels not matching seamlessly. As illustrated in Fig. 1, the boundaries of the MODIS pixels do not perfectly match those of the smaller CDL pixels. Consequently, many CDL pixels located at the boundaries of the MODIS pixels are only partially included within the MODIS pixels. When we count these CDL pixels to calculate the corn ratio, it leads to inaccuracies. Therefore, we use attention mechanism in our model to automatically learn attention, as the weights automatically derived by the attention method will yield more accurate and reliable predictions.
Second, our attention module can automatically learn a general weight for mixed pixels among different datasets. In Fig. 1 and the above experiments, we only use the MODIS dataset to illustrate the cause of mixed pixels and calculate the corn ratio. However, the problem of mixed pixels arises with any feature dataset with a resolution greater than $30\mathrm{m}\times 30\mathrm{m}$. When using many datasets with different resolutions, it would be necessary to calculate different corn ratios for each type of feature and then combine them for prediction, which is not a feasible or elegant solution. Our attention method does not require manually calculating the corn ratio for all datasets; instead, it automatically learns unified weights from different features. This method is more straightforward and elegant. These two reasons also help explain why the correlation between attention and the corn ratio is not a perfect 100% (Sec. 5.3.3), as our attention score is more accurate.
5.6 Limitation of our method
Although our Att-MIL method has been successful in pixel-level corn yield prediction and can effectively address the “mixed pixel” problem, there are still some limitations that need to be considered. First, due to the big size of satellite imagery, we can only extract VIs from it and use these as inputs for the model. This process results in some information loss, which is not entirely consistent with our original goal of pixel-level yield prediction. However, using the complete remote sensing images as input would demand substantial computational resources, making it challenging to balance these two aspects. Second, our method currently uses 16-day average features as input, which means that some subtle environmental changes cannot be captured. However, we consider that a half-month interval is sufficient to capture most relevant features when designing our experiment, and adding more data points would significantly increase the training cost of the model. Therefore, we believe using 16-day intervals is reasonable choice. Finally, while our model performs well in end-of-in-season prediction, it shows weaker performance in the early stages of in-season yield prediction. This reflects the shortcomings of learning-based methods, which require a large amount of data to achieve good predictive performance. In the future, we plan to incorporate process-based models to ensure better performance at earlier stages of corn growth.
6 Conclusion
This paper explores county-level corn yield prediction by treating each county as a collection of cornfield pixels. To address the issues encountered in this specific scenario, we use a combination of MIL and attention mechanism as a solution. First, MIL is employed to leverage the pixel-level RS observations, balance the conflict between computational resources and information integrity, and address the lack of field-level yield records for pixel-level data processing. Subsequently, to solve the mixed pixel problem caused by inconsistent resolutions of feature datasets and crop mask, an attention mechanism is incorporated to assign weights to pixels, thereby enhancing prediction accuracy. We analyze the role of attention from both statistical and micro perspectives, confirming the validity of our approach. In the future, we will try to enhance the generalization of the model by transfer learning in larger-scale corn yield prediction research.
Acknowledgements
We’d like to appreciate the editors and reviewers, whose comments are all valuable and very helpful for improving our manuscript.
Disclosure statement
The authors report no conflict of interest.
Funding
This work was supported by the United States Department of Agriculture (USDA) National Institute of Food and Agriculture, Agriculture and Food Research Initiative Project under Grant 1028199.